OpenStudy (anonymous):

ln(x-3)=ln(5x)+ln(2)

OpenStudy (jdoe0001):

\(\bf ln(x-3)=ln(5x)+ln(2)\\ \quad \\ \textit{recall that}\quad log(a)+log(b)\implies log(a\times b)\\ \quad \\ ln(x-3)=ln(5x)+ln(2)\implies ln(x-3)=ln(5x\times 2)\)